OpenStudy (anonymous):

cos^4t - sin^4t = 1-2sin^2t

OpenStudy (kc_kennylau):

Note that \(\cos^4t-\sin^4t=(\cos^2t-\sin^2t)(\cos^2t+\sin^2t)=\cos^2t-\sin^2t\) :)

OpenStudy (kc_kennylau):

@hartnn I thought you were answering lol

hartnn (hartnn):

i was but i knew you would answer it correctly so i thought i was not needed anymore :)

OpenStudy (kc_kennylau):

oh thanks :)

OpenStudy (kc_kennylau):

And then add \(\sin^2t\) and subtract it again, so that makes: \[\cos^2t+(\sin^2t-\sin^2t)-\sin^2t\]

OpenStudy (kc_kennylau):

Add the \(\cos^2t+\sin^2t\) together to make \(1\) :)

OpenStudy (kc_kennylau):

Complete proof: \[\cos^4t-\sin^4t\]\[=(\cos^2t-\sin^2t)(\cos^2t+\sin^2t)\]\[=\cos^2t-\sin^2t\]\[=\cos^2t+\sin^2t-\sin^2t-\sin^2t\]\[=1-\sin^2t-\sin^2t\]\[=1-2\sin^2t\] Hope I helped you :)

OpenStudy (anonymous):

Thank you!